OpenStudy (anonymous):

Find the slope of the line that goes through the given points. (-2, 0), (0, -3)

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

(-3-0) / (0-(-2)) = (-3) / (2) = -3/2

OpenStudy (anonymous):

0-(-3)/-2-0 =3/-2

OpenStudy (anonymous):

slope = rise/run rise= y1-y2 run= x1-x2